nonlinear analogues of e^{i(kx−ωt)} for integrable systems

Mikhail Kovlayov university of alberta

The similarity between the inverse scattering method for solving nonlinear integrable systems and the Fourier transform method for solving linear PDEs has been known for a while. But do there exist nonlinear anlogues of e^{i(kx−ωt)}, the Fourier representation formula, amplitude modulation and other things naturally appearing as applications of the Fourier transform? The answer will be discussed in the talk.